Technician accidently made 0,150L solution in which sucrose concentration was 3,00 mol/L. Suggest how we could solve this problem from already present solution to create new one with procentive concetration 55% sucrose and aqueous solution mass 100g?

Thank you
1
Expert's answer
2020-09-21T06:34:08-0400

Solution.

"w = \frac{m(sucrose)*100 \%}{m(solution)}"

"m(sucrose) = \frac{w \times m(solution)}{100 \%}"

m(sucrose) = 55 g

"C = \frac{m(sucrose)}{M(sucrose) \times V(solution)}"

C = 1.61 M

"C1 \times V1 = C2 \times V2"

"3.00 \times V1 = 1.61*0.1"

V1 = 0.0537 L = 53.7 ml

So we have to take 53.7 ml of the solution that the technician prepared and dilute it to 100 ml.

Answer:

V1 = 0.0537 L = 53.7 ml

So we have to take 53.7 ml of the solution that the technician prepared and dilute it to 100 ml.